OpenStudy (pokemon23):

The larger of two numbers is 5 more than the smaller.The smaller number plus twice the larger equals 100.Find the numbers. The larger of two numbers is 5 more than the smaller.The smaller number plus twice the larger equals 100.Find the numbers. @Mathematics

OpenStudy (pokemon23):

Saifoo again :)

OpenStudy (saifoo.khan):

the larger of two numbers is 5 more than the smaller. 1) x = y + 5 the smaller number plue twice the larger number equals 100 2) y + 2x = 100 y + 2(y + 5) = 100 {substitute for x from equation 1} y + 2y + 10 = 100 3y = 90 y = 30 1) x = 30 + 5 = 35 .

OpenStudy (saifoo.khan):

[=

OpenStudy (pokemon23):

what is the smaller and larger number?

OpenStudy (pokemon23):

saifoo help